OpenStudy (anonymous):

Please help C(x)=6/x2-9 Find the domain of the rational function type your answer in interval notation

Directrix (directrix):

y= 6/(x^2 - 9) ------------- Zero cannot be the denominator so any value of x that makes x^2 -9 = 0 is excluded from the domain. x^2 -9 = (x +3)(x-3) so x = -3 and x = 3 are excluded from the domain. Domain: (-00,-3) U (-3, 3) U (3, 00)